OpenStudy (anonymous):

What is the approximate distance from (2, 5) to (-2, 1)? 6 Units 5 Units 4 Units 3 Units

OpenStudy (anonymous):

Have you tried using the distance formula?

rishavraj (rishavraj):

distance between two points (x1 , y1) and (x2 , y2) is \[\sqrt{(x_1 -x_2)^2 + (y_1 - y_2)^2}\]

rishavraj (rishavraj):

thts the distance formula :))

OpenStudy (anonymous):

\(\huge \ \sqrt{(2 -(-2))^2 + (5 - 1)^2}\) \(\huge \ \sqrt{(2 +2)^2 + (4)^2}\) \(\huge \ \sqrt{(4)^2 + 16}\) \(\huge \ \sqrt{(16+ 16}\) \(\huge \ \sqrt{(32}\)

OpenStudy (anonymous):

5 units

OpenStudy (anonymous):

^ Wrong.

OpenStudy (anonymous):

it would not be 6

OpenStudy (anonymous):

square 32

OpenStudy (anonymous):

o wait 6 units

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

sorry dump move LOLZ

OpenStudy (anonymous):

*Durp move

OpenStudy (anonymous):

\[\sqrt{32}\approx 5.65685\]

OpenStudy (anonymous):

no see Silly mistake I know its 6 because 6x6+36 5x5=25 whats closer to 32

OpenStudy (anonymous):

see i just rounded down for some really odd reason